What is P after execution? Assume that Z is an integer greater than 2. (solve without selecting actual values, use Z in final answer?

INPUT Z

P=3

DOFOR= 2, Z

P= P+K

ENDDO
Answer #1

The final va;ue of P will be

P=P+K(Z-2)

as the loop will run for the value 2 to Z

Hence every time we are adding K

so P= P+K+K+K....Z-2 times

Hence P=P+K(Z-2)